OpenStudy (anonymous):

I'm trying to find out how to the answer to this equation by these steps. I can't figure out the last few steps. x^2 - 2x - 13= 0 Steps: a) move the constant term to the right side of the equation b) Multiply each term in the equation by four times the coefficient of the x ^2 term. c) square the coefficient of the original x term and add it to both sides of the equation d) take the square root of both sides e) set the left side of the equation to the positive square root of the number on the right side and solve for x f) set the left side of the equation equal to the negative square root of t

OpenStudy (anonymous):

the number on the right side of the equation and solve for x

OpenStudy (anonymous):

Restlt s after step a) x^2 - 2x = 13 b) 4x^2 - 4x = 52 c) 4x^2 - 8x + 4 = 56 d) 2x - 2 = 2*sqrt(14) a) 2x - 2 = -2*sqrt(14)

OpenStudy (anonymous):

use quadratic equation x=-b+/-root(b^2 -4ac) whole divided by 2a.... then sove for x ... you will get the answer x=1+/-root 14

OpenStudy (anonymous):

abtrehearn- would that be -13 instead of 13 because of the first equation

OpenStudy (anonymous):

This problem illustrates an alternative method to the usual complete-the-square nethod.

OpenStudy (anonymous):

The usual method calls for a leading coefficient of 1. The alternative method calls for a perfect square as the leading coefficient.

OpenStudy (anonymous):

Both methods seek a perfect trinomial square for the left side of the equation.